Presentation is loading. Please wait.

Presentation is loading. Please wait.

1 MICE phase I - Project Status Richard Apsimon (Project Manager) & Andy Nichols (Technical Coordinator) … with considerable help from Ken Long! MICE Video.

Similar presentations

Presentation on theme: "1 MICE phase I - Project Status Richard Apsimon (Project Manager) & Andy Nichols (Technical Coordinator) … with considerable help from Ken Long! MICE Video."— Presentation transcript:

1 1 MICE phase I - Project Status Richard Apsimon (Project Manager) & Andy Nichols (Technical Coordinator) … with considerable help from Ken Long! MICE Video Conference 1 st Nov 2007

2 2 Contents 1.MICE WBS 2.MICE target news 3.Upstream beamline 4.DSA & the Decay Solenoid 5.Linde refrigeration installation & commissioning 6.Q35 Refurbishment, testing & installation 7.MICE LCR 8.Other issues


4 4 Weekly MICE Hall planning meetings. Essentially a 2 week look-ahead Weekly MICE Schedule planning meetings, concentrating On the longer term plus planning for the forthcoming weekend

5 5 MICE Target Past problems with bearing wear. Now using DLC (Diamond Like Carbon or Diamolith). It is a “Plasma Enhanced Chemical Vapour Deposition” (PECVD) process that deposits a material layer that is about 3-5um thick. At present only got one, top bearing, DLC coated and the target shaft. >2M operations without wear

6 6 MICE Target Problems that we needed to solve for target installation into ISIS: –Removal of the old target. Need to demonstrate procedure Can we use short Nov. shutdown? –Approval to put the new target into ISIS. YES! MICE procedure accepted by ISIS –Some conditions –Infrastructure ready for target in ISIS. Two targets needed, test vessel, remote handling, fibre optics, control cables. Bottom bearings need coating

7 7 Target: schedule:

8 8 Target: schedule continued: Now planned with DL controls group: needs to be updated.

9 9 Upstream beamline Require to complete/install: –Alignment PROBLEM – need to lower magnet stands –Vacuum vessels PROBLEM with Dipole 1 –Need to move vacuum box –Services: Vacuum Water Electrical –Agreement with M.Krendler (ISIS): –Milestone (deadline) for completion of above 20Jan08: the end of the Christmas shutdown Working with individuals to complete as much as possible before ‘minor maintenance’ period –Next three/four weeks Next opportunity November ‘minor maintenance’ –But most people at the limit of their annual dose. We should get the supplies connected.

10 10 Upstream beam line: power supplies: Rack being installed now: –Required removal of ‘treads’ of north-east stair case Power-supply units and cabling to follow

11 11 Upstream beam line: Xmas work

12 12 DSA (Decay Solenoid Area) June: Hole to synchrotron vault, North wall in place October: Shielding to synchrotron, W & S walls of iron & concrete built. PPS doors installed. Plinth built, additional shielding on plinth, wall painted. Dipole 2 installed, downstream Q35 baseplate in position.

13 13 View over DSA into MICE Hall June: Most of Hall cleared, some shielding October: Decay Solenoid Area civil engineering almost complete. Roof still to be designed & fitted. Can we remove North Wall? On-going discussions. Looks possible if we can replace sandbags with steel shot. More news from Willie & Andy?

14 14 Delay Solenoid So far, major leak has been found and repaired BUT at the cost of considerable damage to MLI. Need 2? weeks of skilled effort. Matt is developing a plan to complete this task. Cold mass still to be aligned and tie rods re-tensioned

15 15 Linde cryogenics After DSA completed, support installed for buffer tank, below which is control plant. Transfer line carries cold helium gas to and from either ‘Test Cap’ or decay solenoid

16 16 Linde cryogenics Helium transfer line installed and connected to ‘Test Cap’. Most of the scaffolding now removed. After cool-down and other tests, transfer line will be connected to PSI Decay Solenoid Installation completed ahead of schedule Commissioning started and hope to finish 25 th or even 18 th Nov.

17 17 Linde Schedule as of 9 th Oct

18 18 Q35 Refurbishment & Installation First 3 magnets finished this week. Test in R6. One technician has left. Replacement being sought. Need to increase # people with crane licences –Trial assembly of magnet support on baseplate complete –Upstream baseplate (Q4-Q6) in DSA has been aligned & levelled (grouted today?) –Downstream plate ready with all components

19 19 Down-stream beam-line: Quadrupole triplets: –First quad triplet refurbished: Magnetic testing to start next week –Trial build of quad stand complete Quadrupole triplets cont.: –Refurbishment of remaining quads: Goal: complete by 20Jan08 –Services being installed now

20 20 Q35 refurbishment: schedule: Focus on Q35 work has paid off, but: –Magnetic testing will start 01Nov07: Developing schedule with Paul Flower now he has returned from CERN –Have had one more coil failure: potential issue: Have expedited testing of remaining coils Continuing refurbishment –Need to consider options when it is known if any more coils have failed

21 21 MICE Local Control Room (MLCR) Rack bases delivered DPM & floor screed completed Network rack delivered this week? PPS rack, populated and tested (@ DL) also delivered this week? Remaining racks the following week. Air conditioning and power to start soon. –Estimated to take three weeks so complete by 9th November. Roof has been fitted & plastered this week Network connection cabling installed. –Connection to RAL site to start as soon as the racks have been delivered from DL. –A wireless access point to be set up as soon as possible. Further wiring into the hall to follow. –Too much civil still going on to start too early.

22 22 MLCR timescale

23 23 Reviews: beam line

24 24 Spectrometer and instrumentation: Spectrometer solenoid: –First spectrometer solenoid: Coils almost complete Expect delivery to FNAL for magnetic field mapping in December –Second spectrometer solenoid: Fabrication advanced Expect delivery ~2 months after first CKov: –Two modules: Most of the first (CKova) at RAL Second module under construction in Mississippi TOF: –TOF0 and TOF1 under construction in Milan: Expected at RAL 15 th November for cosmic ray tests Calorimeter: –Pre-sampler (KL): under construction in Rome Expected at RAL 15 th November for cosmic ray tests –Sandwich calorimeter: Design complete, prototype under construction

25 25 First full tracker being manufactured: –Final step in QA chain (station acceptance rig) now in full production mode –Space frame being built –Support frame in manufacture Electronics: –Being prepared for shipping in FNAL –Will be shipped this week or next week Spectrometer and instrumentation: continued

26 26 Cosmic-ray integration test: Will start 12Nov07: –Will include: CKov, Tracker, ToF, and KL (pre- sampler)

27 27 Other Issues Xmas shutdown –Well planned, WBS manager (Tim Hayler) –Already VERY tight. Every other task that is delayed now moves into this schedule. False floor and plinth in main hall –False floor should be completed in time for decay solenoid to move for cool-down tests Other news –Water chiller order placed. –Rolling platform FDR completed, PRR due. –Magnetic shield. South wall drgs. complete. North wall, design almost complete. Aim to have contracts placed well before end of this FY. –Quotes for “Cyclotron steel” being sought

28 28 Phase I instrumentation: Cosmic integration test: will start 12Nov07 –Craig MacWaters coordinating test Will use R8: –3-phase, water, etc. being planned »Unforeseen expense (likely small, but being estimated) –Components: Tracker: –Station acceptance QA rig now running at Imperial »Station ‘5’ (first final station) has been QAed: results OK; »Decide today 15:30(?): proceed with QA or more studies required on station 5 –Space frame in manufacture –Cosmic stand signed off for manufacture (?) TOF: –Expected to arrive 15Nov07 KL: –Expected to arrive 15Nov07 CKova –Already at RAL, needs to be assembled DAQ –Already at RAL, needs to be moved to R8

29 29 Cooling channel components: RFCC module: –CC design and fabrication done in collaboration by LBNL and ICST in Harbin, China conductor ready to order –RF cavities will be similar to existing MuCool prototype fabrication to get under way shortly –MuCool 201 MHz cavity R&D: Without field now have 19 MV/M

30 30 Cooling channel components: AFC module: Focus coil module presently out for tender –anticipate contract award in September 2007 –two coils that can run with same or opposite polarity »20-L LH 2 absorber (plus safety windows) fits inside Oxford, RAL, KEK

31 31 Current concerns Preparation of –The PSI solenoid –MICE target –Upstream beamline Is the misalignment critical? Can it be solved in time? –Connection of upstream & downstream beamline components Radiation surveys –Many areas where these matter; Target removal Connection of magnets in synchrotron vault Work inside DSA – removal of North wall Excavation of earth for H2 vents and pumps

Download ppt "1 MICE phase I - Project Status Richard Apsimon (Project Manager) & Andy Nichols (Technical Coordinator) … with considerable help from Ken Long! MICE Video."

Similar presentations

Ads by Google